Find the area of an equilateral triangle given a side length of 8 in.
Answer by unlockmath(1688)

Hello,
First, let's find the height by using:
A^2+4^2=8^2
Rewritten as:
A^2=48
A=sq rt 48
Now we can use the area of a triangle:
1/2*4*sq rt 48 to get:
2 sq rt 48 or rewritten as:
4 sq rt 12 square inches
make sense?
RJ
www.math-unlock.com